Question Help In a certain country, the true probability of a baty being a of them is a girt? boy is 0519. Among the next sik randomy selected births in the country, wtprobability that at least one The probability is (Round to three decimal places as needed)
Answer #1

The true probability of baby being boy = 0.519, So the true probability that a baby being girl is = 1 - 0.519 = 0.481.

We have 6 randomly selected birth

Let X be the number of baby being girl in these 6 births.

P( X is at least 1 ) = P( X \geq 1 ) = 1 - P( none of baby being girl)

                                                = 1 - P( X = 0 )

                                                = 1 - ( 0.519)^6

                                                = 1 - 0.020

Probability is                         = 0.980